G-081 field Season 2012-2013

Logbook entries (add new)

Date Title Preview People
Jan. 8, 2013 Attempt to set up live pan-tilt netcam in Heroine tower After dinner (21:00) went to Heroine alone to set up pan-tilt. Solar controller ... Aaron Curtis, edit
Jan. 8, 2013 Derigging Mammoth with copter test After dinner (9pm), went to Heroine alone to set up pan-tilt. Solar controller w... Aaron Curtis, edit
Jan. 5, 2013 3D DTS pullout Double-bagged and removed 7 discolored bamboos & surrounding soil Ice bucket ... Tim Burton, Aaron Curtis, Nial Peters, edit
Jan. 4, 2013 Cinemax DTS cable distance calibration heating 1L Boiling H20 10:54 NZDT @ 1924-1940ft 1.5L Boiling 11:05 NZDT @ 1996-2004m no... Sune Tamm, Tim Burton, Aaron Curtis, edit
Dec. 21, 2012 Tinytag downloading Managed to download tinytags. 524764: Cathedral top 524754: Cathedral bottom ... Aaron Curtis, edit
Dec. 20, 2012 [DATE ESTIMATED] Retrieval of HOBOs from back of Writing this from memory in Socorro at 14 May. Not sure about the date. This was... Tim Burton, Sune Tamm, Aaron Curtis, edit
Dec. 18, 2012 Cathedral DTS cable removal Cathedral DTS cable removal. Notebook has drawing. Aaron Curtis, Tim Burton, edit
Dec. 15, 2012 Lidar, splicing Down mammoth with Clive, Drea, Jed. Clive filming, helping lidar crew. Me failin... Clive Oppenheimer, Jed Frechette, Drea Killingsworth, Aaron Curtis, edit
Dec. 12, 2012 Datalogger downloading At LEH, downloading dataloggers that had recently been retrieved from caves 5... Aaron Curtis, edit
Dec. 10, 2012 3D DTS installation completed Sune, Tim, Aaron left LEH 9:30, down Mammoth. Set up DTS 3D web in Cinemax (see ... Sune Tamm, Tim Burton, Aaron Curtis, edit
Dec. 9, 2012 Adding tarp to melt slush bath Went back down Mammoth after dinner around 2130. Quick tour, put tarp over bucke... Nial Peters, Jed Frechette, Aaron Curtis, edit
Dec. 9, 2012 Gas pulse in Warren? Helped Nial w/ rim power cable in the morning. When returned to LEH, Drea & J... Toby Fischer, Aaron Curtis, edit
Dec. 8, 2012 Beginning of 3D DTS installation Underground about 11am in Mammoth. Checked other end of Cathedral cave, no light... Sune Tamm, Tim Burton, Aaron Curtis, edit
Dec. 7, 2012 Warren gas sampling Pulled out 2 sampling systems in middle passage (E-W) past V2. Set up systems in... Toby Fischer, Aaron Curtis, edit
Dec. 6, 2012 Gas sampling in Warren Down warren w/ Toby. Put 3x vent sampling apparati in place. 1 in V2, one in V5 ... Toby Fischer, Aaron Curtis, edit
Dec. 5, 2012 Hut day Storm. Got webcam and FLIR on website. Aaron Curtis, edit
Dec. 4, 2012 Warren east passage datalogger removal Underground by 2pm. Showed Toby V2, VHub [hard to read-- implies "hubert's vent"... Toby Fischer, Aaron Curtis, edit
Dec. 3, 2012 Ice tower ridge datalogger maintenance Left for ITR (Tim & Aaron) @ 9am. Went in Heroine. Most batteries < 1V, few 1 to... Tim Burton, Aaron Curtis, edit
Dec. 3, 2012 Ice tower ridge datalogger maintenance Left for ITR (Tim & Aaron) @ 9am. Went in Heroine. Most batteries < 1V, few 1 to...   edit
Dec. 3, 2012 Ice tower ridge datalogger maintenance Left for ITR (Tim & Aaron) @ 9am. Went in Heroine. Most batteries < 1V, few 1 to...   edit
Dec. 3, 2012 Ice tower ridge datalogger maintenance Left for ITR (Tim & Aaron) @ 9am. Went in Heroine. Most batteries < 1V, few 1 to...   edit
Dec. 2, 2012 DTS situation check and first entry into Cathedral of the season Aaron & Tim went down Mammoth. Remove datalogger, sensors (incl smashed ultrason... Aaron Curtis, Tim Burton, edit
Nov. 30, 2012 Mammoth rigging Left for Kachina cave. (w/ tim?) Went in Kachina downhill. Pulled datalogger 239... Tim Burton, Aaron Curtis, edit
Nov. 29, 2012 Crevasse Cave Established entry to cave Tim Burton, Alasdair Turner, Sune Tamm, edit
Nov. 29, 2012 Warren second FTIR trip of season Solomon, Clive, Aaron underground @ 1430. Lugged FTIR Gear to V5. Started Agata ...   edit
Nov. 28, 2012 rigging trip and exploration pre biology second visit to the cave (first in 08/09) easy rappel this time down a few metre... Tim Burton, Alasdair Turner, edit
Nov. 28, 2012 Kachina Cave Established entry to cave Tim Burton, Alasdair Turner, edit
Nov. 28, 2012 Wormtongue Hole Established entry to cave and verified good location for samples. Tim Burton, Alasdair Turner, edit
Nov. 28, 2012 Warren FTIR Cave FTIR Mission with Clive and Nial. Put gas box "Agata" in tube structure und... Clive Oppenheimer, Aaron Curtis, Nial Peters, edit
Nov. 27, 2012 First Descent Two soil samples taken for biological monitoring. Alasdair Turner, Hubert Staudigel, edit
Nov. 27, 2012 Warren entrance FTIR Clive, Aaron did FTIR on Warren ent, while Brad, Hubert, Tony collected more sam... Clive Oppenheimer, Aaron Curtis, Bradley Tebo, Hubert Staudigel, Anthony Rigoni, edit
Nov. 27, 2012 Initial Mammoth trip 16:00 departure. Gaetano, Aaron, Tim, Alasdair rigged Mammoth same ent as last y... Gaetano Giudice, Aaron Curtis, Tim Burton, Alasdair Turner, edit
Nov. 26, 2012 Warren Cave Rappelled into Warren cave. Collected G-439 experiments and sampled soil, rock, ... Aaron Curtis, Tim Burton, Alasdair Turner, Bradley Tebo, Anthony Rigoni, Hubert Staudigel, edit
Nov. 26, 2012 Datalogger system removal from Warren Brad, hubert, tony, alasdair, tim, aaron to warren. Got going around 9:10. Found... Bradley Tebo, Hubert Staudigel, Anthony Rigoni, Tim Burton, Aaron Curtis, edit
Nov. 25, 2012 Warren Cave Rappelled in at main vent, proceeded to V2 (vent 2) area and began exploring new... Alasdair Turner, Anthony Rigoni, Aaron Curtis, Sune Tamm, edit
Nov. 25, 2012 gaining access to warren - Drove up to LEH, arrived 8am 10am - Tony, Alasdair, Sune, Aaron --> Warren ... Bradley Tebo, Alasdair Turner, Anthony Rigoni, Hubert Staudigel, Aaron Curtis, edit
loading page...